# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v2

@@ Abstract
Continual learning (CL) with large pre-trained models is challenged by catastrophic forgetting and task interference. Existing LoRA-based Mixture-of-Experts (MoE) approaches mitigate forgetting by assigning and freezing task-specific adapters, but suffer from interference, redundancy, and ambiguous routing due to coarse adapter-level selection. However, this design introduces three key challenges: 1) Interference: Activating full LoRA experts per input leads to subspace interference and prevents selective reuse of useful components across tasks. 2) Redundancy: Newly added experts often duplicate or contradict existing knowledge due to unnecessary activation of unrelated ranks and insufficient reuse of relevant ones. 3) Ambiguity: Overlapping features across tasks confuse the router, resulting in unstable expert assignments. As more experts accumulate, earlier task routing degrades, accelerating forgetting. We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparse rank activation for CL. Unlike mixing multiple low-rank matrices, MoRA decomposes each rank-r update into r rank-one components, each treated as an independent expert, enabling fine-grained rank-one expert utilization while mitigating interference and redundancy. To avoid ambiguous routing, we propose that each rank-one expert can infer its own relevance via intermediate activations. Coupled with our proposed rank pruning and activation budgets, MoRA adaptively selects a sparse mixture of ranks per input. We validate MoRA on continual learning benchmarks using CLIP and language models, analyzing both in-domain learning and out-of-domain forgetting/generalization during fine-tuning. MoRA shows significant effectiveness in enhancing CL with PTMs, and improving generalization while mitigating forgetting.
@@ 1 Introduction
Continual learning (CL) (Hadsell et al., 2020; De Lange et al., 2021; Ding et al., 2022) aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting (Nguyen et al., 2019; McCloskey & Cohen, 1989) or the need for repeated fine-tuning on all previously seen data (Wang et al., 2022e; 2024). The emergence of large pre-trained models (PTMs), including Vision Transformer (ViT) for vision (Dosovitskiy et al., 2020), Language Model (LM) (Grattafiori et al., 2024; Raffel et al., 2020), and CLIP for vision–language embeddings (Radford et al., 2021), has spurred a wave exploring CL with PTMs (Xu et al., 2024; Wang et al., 2023). Despite the success of PTMs, CL on them across sequential tasks still faces key challenges, including preventing catastrophic forgetting and preserving, or even enhancing, the pre-trained capabilities to generalize to unseen or future cases (Qiao & Mahdavi, 2024).
A common way to integrate new knowledge into PTMs is full fine-tuning (Zheng et al., 2023; Garg et al., 2023) or PEFT with a small set of learnable parameters (Yang et al., 2024b), such as LoRA (Hu et al., 2021) (Fig. 1(a)). However, continually updating the same parameters often leads to interference and forgetting, even with a single new task (Biderman et al., 2024). Leveraging LoRA’s regularization and flexibility, many CL methods adopt LoRA or its variants with additional strategies to mitigate forgetting (Yu et al., 2024; Wang et al., 2023; Qiao & Mahdavi, 2024).
@@
Inspired by mixture-of-experts (MoE) architectures (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024) and the modular nature of LoRA adapters, MoE frameworks using LoRA adapters as experts have been studied (Dou et al., 2023; Wu et al., 2024b) and applied in CL (Fig. 1(b)) have been widely adopted in CL (Yu et al., 2024; Wang et al., 2024; Yang et al., 2024a; Chen et al., 2024; Li et al., 2025). These works either predefine an MoE with LoRA for CL (Yang et al., 2024a), or incrementally add experts (Wang et al., 2024) or task-specific routers (Yu et al., 2024), assuming MoE benefits CL by isolating task interference. Such methods (Wang et al., 2024; Qiao & Mahdavi, 2024; Rusu et al., 2016; Yu et al., 2024) freeze old components and add new ones (e.g., experts or routers) to reduce forgetting. Despite design differences, we collectively refer a plain and general design with LoRA-based MoE as MoE-LoRA. In MoE-LoRA models, each expert is a LoRA adapter with multiple ranks (subspaces), and the router selects among experts with each LoRA adapter as a unit. Despite their success, we argue that the coarse granularity of experts and routing introduces three key challenges that limit effective utilization:
(1) Interference across experts and limited knowledge reuse. A coarse-grained multi-rank LoRA expert is packed with a rich set of knowledge that must be fully activated or not at all. For each sample (i.e., token), any two experts may hold both complementary and conflicting knowledge, creating a dilemma: selecting both causes interference, while only selecting either one prevents reuse.
(2) Learning redundancy and conflicts. As a result of (1), (newly added) experts often have to carry conflicting or redundant information, either from activating other experts with both useful and unnecessary content or from insufficient reuse of existing ones. This coarse selection exacerbates interference and forgetting during inference.
(3) Routing ambiguity and forgetting. With coarse-grained experts and routing, redundancy and conflicts hinder precise input–module assignment, leading to suboptimal mixtures (Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024). As more experts accumulate in CL, router assignments for earlier tasks grow unreliable, increasing forgetting.
@@
The three challenges are intertwined causes and effects of coarse-grained experts and routing, limiting the potential of promising MoE-LoRA design.
Figure 1: Conceptional illustration of CL with (a) LoRA, (b) MoE-LoRA, and (c) MoRA (Ours).
We propose MoRA, a Mixture-of-Rank Adaptive learning approaches with self-activated and sparsely selected ranks across the LoRAs added at different tasks. To address the limitations of coarse-grained MoE-LoRA methods, MoRA operates at a fine-grained level by treating each rank-one subspace within a LoRA adapter as an individual expert (or treating each rank-one LoRA as an expert). Expert selection and routing occur at the rank level, avoiding interference from whole LoRA blocks. This decomposition promotes specialization on narrower data regions, reduces redundancy, and enables effective knowledge reuse across tasks.
We introduce a self-activation mechanism to select a sparse set of ranks without relying on a separately trained router. By defining each rank-one component as an expert, we avoid a separate router for expert assignment: the rank-one $\mathbf{A}$ in LoRA can act as a weighting router for its $\mathbf{B}$. Each rank evaluates its relevance to the input and determines its contribution, grounded in the view of low-rank updates as linear key–value memory. This design unifies rank-one LoRA with fine-grained MoE while preserving the original characteristics, saving parameters (for additional router), reducing routing confusion, and further mitigating forgetting (Fig. 1(c)). Sparsity is enforced via temperature-scaled softmax together with top-k masking and rank pruning under a rank activation budget. As a complement, test-time thresholding further improves adaptability across cases. Leveraging the connection between PTMs and low-dimensional manifolds (Li et al., 2018; Aghajanyan et al., 2020), MoRA enables precise, input-specific adaptation through selective subspace updates. With fine-grained rank-level experts and sparse activation, MoRA enables continual learning by reusing old ranks and selectively activating a few new ones per sample, allowing efficient “little-by-little” adaptation with minimal redundancy. This sample-specific activation at inference also preserves pre-trained knowledge and improves generalization to unseen tasks.
We validate MoRA on CL with both vision–language CLIP and LMs, analyzing its impact on pre-trained knowledge during low-rank fine-tuning. MoRA delivers strong downstream performance with far fewer active parameters, greatly reduces forgetting of prior tasks and pre-trained knowledge, and even improves generalization to unseen domains.
@@ 2 Related Work
Continual learning enables sequential knowledge acquisition without forgetting. Experience replay (ER) methods (Luo et al., 2023; Aljundi et al., 2019b; Chaudhry et al., 2018a; Liu et al., 2020; Chaudhry et al., 2018b; Yan et al., 2022; 2021) interleave past examples with new data. Parameter regularization (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Aljundi et al., 2019a; Jha et al., 2023) penalizes updates to critical weights. Dynamic networks (Zhou et al., 2024; Wang et al., 2022a; b; Zhou et al., 2022; Wang et al., 2024; McDonnell et al., 2024; Liang & Li, 2024; Wang et al., 2022e; d; Smith et al., 2023; Wang et al., 2022c; 2024) allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
Continual learning of PTMs. For CL on vision–language CLIP model (Jha et al., 2024; Zhang et al., 2024b; Garg et al., 2023), methods like ZSCL (Zheng et al., 2023) retain zero-shot performance during adaptation, and follow-up work (Yu et al., 2024; Tang et al., 2025; Xu et al., 2024; Lu et al., 2024) continually fine-tunes while leveraging frozen pre-trained predictions. The X-TAIL benchmark (Xu et al., 2024) further challenges models by mixing domain labels at test time. In language models (LMs) (de Masson D’Autume et al., 2019; Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024), continual learning uses capacity expansion or task-specific submodules to reduce interference.
Low‐rank adaptation (LoRA) (Hu et al., 2021) is widely used for parameter‐efficient fine‐tuning of large pre‐trained models. Building on this foundation, recent methods reformulate LoRA’s updates via SVD‐based initialization and dynamic rank scheduling (Zhang et al., 2024a; Meng et al., 2024; Ding et al., 2023; Zhang et al., 2023; Liu et al., 2024; Wu et al., 2024a), showing that task adaptation primarily lies in a compact subspace. In this work, we offer a complementary perspective by viewing both pre‐trained weight matrix and its low‐rank updates as a linear associative memory (Li et al., 2018; Aghajanyan et al., 2020; Kohonen, 1972; Anderson, 1972). Under this lens, a rank-$r$ update corresponds to writing $r$ new key–value entries into the pre-trained memory matrix, with each rank-one component acting as an independent memory slot.
Mixture‐of‐experts (MoE) in LoRA fine‐tuning. MoE increases model capacity by replacing the Transformer’s feed-forward layer with expert subnetworks, routing each input to a small subset and using load‐balancing losses to even out usage (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Dai et al., 2024). This paradigm has been adapted for standard fine-tuning (Chen et al., 2023; Li et al., 2024; Dou et al., 2023) and CL (Yu et al., 2024; Wang et al., 2024; Yang et al., 2024a; Chen et al., 2024) with LoRA, treating each low‐rank adapter as an expert: new adapters are trained and frozen per task to prevent forgetting. In contrast, we decompose each rank‐$r$ update into $r$ rank‐11 components and compute an input‐dependent mixture over these fine‐grained experts, greatly enhancing both expert specialization and mixture diversity.
@@ 3.1 Preliminaries
Continual learning. In CL, a model sequentially learns $T$ tasks. For task $t\in{1,\dots,T}$, let $\mathcal{D}^{t}={(\mathbf{x}^{t}_{i},y^{t}_{i})}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{n\times d}$, $y^{t}_{i}\in\mathcal{C}^{t}$, and $N^{t}$ is the number of examples. In the memory-free setting, the model may access only $\mathcal{D}^{t}$ and cannot access data from any $\mathcal{D}^{u}$ with $u<t$.
Low-rank adaptation. LoRA (Hu et al., 2021) parameterizes a low‐rank update to a pre‐trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{out}\times d_{in}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{out}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{in},d_{out})$. The updated weight matrix is then defined as:
@@
In this formulation, the original weights $\mathbf{W}_{0}$ remain fixed, and only $\mathbf{B}$ and $\mathbf{A}$ are trained, reducing the number of trainable parameters from $d_{in}d_{out}$ to $r(d_{in}+d_{out})$.
@@
Mixture‐of‐Experts LoRA. Building on the Mixture‐of‐Experts (MoE) paradigm, a common generic framework of MoE‐LoRA (Yu et al., 2024; Wang et al., 2024) treats each LoRA as an independent expert. Suppose after $T$ tasks, we have $T$ LoRA experts $\{(\mathbf{A}^{1},\mathbf{B}^{1}),\dots,(\mathbf{A}^{T},\mathbf{B}^{T})\}$. For an input token $x\in\mathbb{R}^{d_{\rm in}}$, the overall LoRA update in this framework is given by $\Delta\mathbf{W}=\sum\nolimits_{i=1}^{T}R(x)_{i}\mathbf{B}^{i}\,\mathbf{A}^{i},$ where the mixture weight $R(x)\in\mathbb{R}^{T}$ is produced by a learned router $R(\cdot)=\mathrm{softmax}(x\mathbf{W}_{r})$ and $\mathbf{W}_{r}\in\mathbb{R}^{d_{\rm in}\times T}$ contains the router’s trainable parameters. Each LoRA’s contribution is weighted by the learnable router, enabling the model to dynamically select and combine the most relevant low‐rank updates for each token. In practice, a Top-$k$ masking is typically applied to mixture weights to enforce sparsity, activating only the $k$ most relevant experts and controlling computational overhead.
3.2 Model Weights and Updates as Linear Associative Memory Model
Weight matrix as linear associative memory. Large pre‐trained models (PTMs) have been shown to inhabit a low “intrinsic‐dimension” manifold compared to their full parameter count (Li et al., 2018; Aghajanyan et al., 2020). Any weight matrix of a pre-trained model $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ can be treated as a linear associative memory model with a key-value structure (Kohonen, 1972; Anderson, 1972; Bau et al., 2020; Meng et al., 2022). Concretely, for any matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ and input $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$, the product $\mathbf{W}\,\mathbf{x}$ implements a content‐addressable memory read over $m$ stored key–value pairs $\{(\mathbf{K}_{i,:},\mathbf{V}_{:,i})\}_{i=1}^{m}$, with $m$ potentially large and varied across each weight matrix. Here, ${\mathbf{K}}=\begin{bmatrix}{\mathbf{K}}_{1,:}^{\top}&\cdots&{\mathbf{K}}_{m,:}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m\times d_{\text{in}}}$ and $\mathbf{V}$$\begin{bmatrix}{\mathbf{V}}_{:,1}&\cdots&{\mathbf{V}}_{:,m}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times m}$ with $\mathbf{K}_{i,:}\in\mathbb{R}^{1\times d_{\text{in}}}$ is the $i$-th row of $\mathbf{K}$} and ${\mathbf{V}}_{:,i}\in\mathbb{R}^{d_{\text{out}}\times 1}$ is the $i$-th column of $\mathbf{V}$}. Given any input $\mathbf{x}$, the output $\mathbf{y}\in\mathbb{R}^{d_{\text{out}}}$ is computed as
$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{V}_{:,i}(\mathbf{K}_{i,:}\mathbf{x}).$ (2)
In this formulation, the forward pass through the weight matrix with the input decomposes into two steps: (1) computing response scores or assignment weights via the key vectors $\mathbf{K}_{i,:}$, and (2) aggregating the corresponding value vectors $\mathbf{V}_{:,i}$ according to those weights.

### v3

# Little By Little: Continual Learning via Incremental Mixture of Rank-1 Associative Memory Experts

@@ Abstract
Continual learning (CL) with large pre-trained models is challenged by task interference and catastrophic forgetting. Existing LoRA-based Mixture-of-Experts (MoE) methods mitigate forgetting by adding new task-specific adapters and freezing old ones, but often suffer from redundancy, interference, and ambiguous routing due to coarse-grained experts and routing. Coarse-grained experts (i.e., full LoRA adapters with large rank) encode low-specialty information. Newly added experts often duplicate or conflict with existing ones, causing redundancy and interference. Their low specialization further confuses the router, accelerating routing degradation and forgetting as experts accumulate. In this work, we propose MoRAM (Mixture of Rank-1 Associative Memory). Grounded in the view that weight matrices function as linear associative memories, MoRAM achieves CL as gradual incrementing of atomic rank-1 memory experts. Each rank-1 adapter acts as a fine-grained MoE expert or an associative memory unit. By viewing rank-1 adapters as key–value pairs, we eliminate explicit routers in MoE-LoRA, using a self-activation mechanism where each memory atom evaluates its own relevance via its intrinsic key. This transforms the adaptation process into robust, content-addressable retrieval. Extensive experiments on CLIP and LLMs demonstrate that MoRAM significantly outperforms state-of-the-art baselines, achieving superior plasticity-stability trade-offs, improving generalization while mitigating forgetting.
@@ 1 Introduction
Continual learning (CL) [29, 16, 20] aims to enable models to incrementally and efficiently acquire new knowledge from a stream of tasks, without catastrophic forgetting [57, 54] or the need for repeated fine-tuning on all previously seen data [85, 78]. With the ascendancy of Large Pre-trained Models (PTMs) in vision [21, 63] and language [64, 28], the challenge has shifted: how do we efficiently insert specific new capabilities into a massive, static memory structure without disrupting its existing generalized knowledge?
Figure 1: Conceptual illustration of CL with (a) LoRA, (b) MoE-LoRA, and (c) MoRAM (Ours).
@@
A dominant approach for adapting PTMs is Parameter-Efficient Fine-Tuning (PEFT), particularly Low-Rank Adaptation (LoRA, Fig. 1(a)) [33]. While efficient, standard LoRA applies updates as dense modifications to the weight space. This creates a fundamental conflict: to learn a new task, the model must inevitably alter the representation space used by previous tasks, leading to interference and forgetting [6]. To mitigate this, recent works have adopted Mixture-of-Experts MoE frameworks using LoRA adapters as experts (Fig. 1(b)) have been studied [22, 88] and have been widely adopted in CL [96, 78, 94, 12, 43]. These works either predefine an MoE with LoRA for CL [94], or incrementally add experts [78] or task-specific routers [96], assuming MoE benefits CL by isolating task interference. Such methods [78, 61, 67, 96] freeze old components and add new ones (e.g., experts or routers) to reduce forgetting. Despite design differences, we collectively refer to a plain and general design with LoRA-based MoE as MoE-LoRA. In MoE-LoRA models, each expert is a LoRA adapter with multiple ranks (in subspaces), and the router selects among experts with each LoRA adapter as a unit.
However, current MoE-LoRA approaches operate at a coarse granularity that limits their effectiveness in CL. Treating entire rank-$r$ adapters as indivisible experts forces each coarse-grained expert to be activated and learned as a single unit. It enables each expert to learn a broad range of information, leading to low specialty and combinatorial expressivity [30, 51, 15]. It specifically introduces three challenges:
(1) Interference: A coarse expert contains a mix of low-specialty knowledge. Activating coarse-grained experts (i.e., LoRA adapters with large rank) for a specific input inevitably triggers irrelevant subspaces, causing interference.
(2) Redundancy: New experts cannot selectively reuse specific atoms of knowledge from old experts; if the limited combination cannot cover the new task, they must relearn entire blocks, leading to inefficient capacity usage.
(3) Routing/Retrieval Collapse: The low-specialization of coarse-grained experts further confuse the router. As the number of experts grows, routers increasingly struggle to reliably index the expanding expert pool (i.e., “routing ambiguity”), leading to routing collapse and drift of old experts, and ultimately accelerating forgetting.
@@
The coarse-grained experts and these resulting challenges limit the potential of the promising MoE-LoRA design.
To address the limitations, we fundamentally rethink adaptation by synthesizing the intrinsic dimensionality of PTMs [1] through the lens of Linear Associative Memory [37]. While PTMs reside in a high-dimensional parameter space, they operate on a low-rank manifold, effectively functioning as a composition of atomic key-value pairs of a memory system [37]. From this perspective, the optimal CL update is not a “dense matrix” or “coarse adapters”, but an associative memory system with expandable and retrievable atomic memory units. Conventional methods that update entire matrices of PTM inevitably alter the global interference patterns of the memory, easily causing catastrophic forgetting. Coarse-grained MoE-LoRA methods suffer from brittle expert expansion and retrieval. From a memory-augmentation perspective, we treat CL as incremental external memory expansion with fine-grained atomic memory units for specialization and efficiency, while casting inference as input-specific memory retrieval and application.
Based on this insight, we propose MoRAM (Mixture of Rank-1 Associative Memory), a framework that continually updates PTMs by incrementing an associative memory system with atomic rank-1 experts “little by little” (Fig. 1(c)). The associative memory can also be seen as a mixture-of-expert model with rank-1 adapters as experts. Each rank-1 adapter acts as a fine-grained MoE expert or an atomic memory unit. By viewing rank-1 experts as key–value memory pairs, we can naturally eliminate the additional explicit routers in MoE-LoRA. With the proposed self-activation mechanism, each expert can evaluate its own relevance to the input via its intrinsic key. This shifts the paradigm from address-based routing (learning where to send data) to content-addressable retrieval (inputs automatically triggering the correct memory). This operation eliminates the source of routing collapse and forgetting from the additional router, increasing the reliability of the memory retrieval, even as the memory bank expands, while reducing computation.
Our contributions are as follows: (1) We introduce MoRAM, a novel CL framework that treats Rank-1 updates as atomic associative memory experts, enabling fine-grained knowledge reuse and minimizing interference. (2) We propose a sparse self-activation routing mechanism that eliminates external routers, allowing the model to sparsely retrieve relevant memory atoms based on input content alone. (3) We conduct extensive evaluations on both vision–language (CLIP) and large language model (LLM) benchmarks. Empirical results demonstrate that MoRAM significantly outperforms state-of-the-art methods, achieving superior plasticity-stability trade-offs, while effectively preserving pre-trained generalization capabilities.
@@ 2 Related Work
Continual learning enables sequential knowledge acquisition without forgetting. Experience replay (ER) methods [52, 4, 9, 46, 10, 92, 93] interleave past examples with new data. Parameter regularization [36, 2, 97, 3, 35] penalizes updates to critical weights. Dynamic networks [106, 76, 77, 107, 78, 55, 45, 85, 84, 69, 82, 78] allocate new capacity on the fly and preserve dedicated pathways for prior tasks.
Continual learning of PTMs. For CL on vision–language CLIP model [34, 101, 27], methods like ZSCL [105] retain zero-shot performance during adaptation, and follow-up work [96, 70, 91, 50, 89] continually fine-tunes while leveraging frozen pre-trained predictions. The X-TAIL benchmark [91] further challenges models by mixing domain labels at test time. In language models (LMs) [17, 62, 65, 80, 61], continual learning uses capacity expansion or task-specific submodules to reduce interference.
Low‐rank adaptation (LoRA) [33] is widely used for parameter-efficient fine-tuning of large pre-trained models. Building on this foundation, recent methods have reformulated LoRA’s updates via SVD-based initialization and dynamic rank scheduling [98, 56, 19, 100, 47, 87], demonstrating that task adaptation primarily relies on finding a compact subspace. In this work, we offer a complementary perspective by treating both the pre-trained weight matrix and its low-rank updates through the lens of linear associative memory [41, 1, 37, 5]. Under this formulation, a rank-$r$ update corresponds to $r$ new memory entries into the matrix, where each rank-1 component is an atomic memory slot.
Mixture-of-Experts (MoE) with LoRA. MoE scales capacity by routing inputs to sparse expert subnetworks via load-balancing [68, 40, 24, 15]. This paradigm has been adapted for LoRA fine-tuning [13, 42, 22] and CL [96, 78, 94, 12] by treating adapters as experts, typically frozen per task to prevent forgetting. In contrast, we decompose rank-$r$ updates into $r$ atomic rank-1 components and compute an input-dependent mixture over these fine-grained experts, significantly enhancing specialization and diversity.
@@ 3.1 Preliminaries
Continual learning. In CL, a model sequentially learns $T$ tasks. For task $t\in{1,\dots,T}$, let $\mathcal{D}^{t}={\{}(\mathbf{x}^{t}_{i},y^{t}_{i}){\}}_{i=1}^{N^{t}}$, where $\mathbf{x}^{t}_{i}\in\mathbb{R}^{n\times d}$, $y^{t}_{i}\in\mathcal{C}^{t}$, and $N^{t}$ is the number of examples. In the memory-free setting, the model may access only $\mathcal{D}^{t}$ and cannot access data from any $\mathcal{D}^{u}$ with $u<t$.
Low-rank adaptation. LoRA [33] parameterizes a low‐rank update to a pre‐trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ by introducing two factors $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{\text{in}}}$, such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$, where $r\ll\min(d_{\text{in}},d_{\text{out}})$. The updated weight matrix is then defined as:
@@
In this formulation, the original weights $\mathbf{W}_{0}$ remain fixed, and only $\mathbf{B}$ and $\mathbf{A}$ are trained, reducing the number of trainable parameters from $d_{\text{in}}d_{\text{out}}$ to $r(d_{\text{in}}+d_{\text{out}})$. While parameter-efficient, standard LoRA applies the rank-$r$ update densely to every input. This would result in updates for a new task that inevitably perturbs the subspaces used by prior tasks, leading to interference.
@@
Mixture‐of‐Experts LoRA. Building on the Mixture‐of‐Experts (MoE) paradigm, a common generic framework of MoE‐LoRA [96, 78] treats each LoRA as an independent expert. Suppose after $T$ tasks, we have $T$ LoRA experts $\{(\mathbf{A}^{1},\mathbf{B}^{1}),\dots,(\mathbf{A}^{T},\mathbf{B}^{T})\}$. For an input token $x\in\mathbb{R}^{d_{\rm in}}$, the overall LoRA update in this framework is given by $\Delta\mathbf{W}=\sum\nolimits_{i=1}^{T}R(x)_{i}\mathbf{B}^{i}\,\mathbf{A}^{i},$ where the mixture weight $R(x)\in\mathbb{R}^{T}$ is produced by a learned router $R(\cdot)=\mathrm{softmax}(x\mathbf{W}_{r})$ and $\mathbf{W}_{r}\in\mathbb{R}^{d_{\rm in}\times T}$ contains the router’s trainable parameters. Each LoRA’s contribution is weighted by the learnable router, enabling the model to dynamically select and combine the most relevant low‐rank updates for each token. In practice, a top-k masking is typically applied to mixture weights to enforce sparsity, activating only the $k$ most relevant experts. However, this design remains coarse-grained. Each expert is a dense rank-$r$ block, and the external router $R(\mathbf{x})$ must learn to map inputs to these overlapping experts. As the number of tasks grows, this external mapping becomes ambiguous, leading to routing collapse and forgetting.